Physics Iceberg question?

The volume of an iceberg is 664,000 ft^3.

(a)
What is its weight (in lb), assuming it is pure ice? (Round your answer to at least four significant figures.)
____ lb

(b)
What is the volume of seawater (in ft^3) it displaces when floating? (Hint: You know what the weight of the seawater is. Round your answer to at least three significant figures.)
____ ft cubed

(c)
What is the volume of the part of the iceberg out of the water (in ft^3)?
____ ft cubed

Thank you for the help!
1
Expert's answer
2018-06-11T15:29:08-0400
(a)The density of ice is ρ=57.23 lb/〖ft〗^3 . Thus, the weight of the V=664000〖ft〗^(3.)is m=ρV=38,000,720lb.
(b) The weight of seawater equals displaced by the iceberg equals to the weight of iceberg. The density of seawater is ρ_w=63.93 lb/〖ft〗^3 . The volume of seawater is V_w=m/ρ_w =594,411.387〖ft〗^3.
(c) the volume of the part of the iceberg out of the water is V-V_w=69,588.613〖ft〗^3.